# Spectral Moment Formulae for $\hbox{GL}(3)\times\hbox{GL}(2)$ L-functions II: The Eisenstein Case

### v2

@@ 1.4.2. Main terms
Recent works on Moment Conjectures (e.g., [Pe15] and [CIS12]) have shown that it is both technically and aesthetically preferable to state and prove results in terms of complete (rather than incomplete) $L$-functions. Indeed, the complete set of the main terms of a moment of $L$-functions is expected to reflect the symmetries originating from successive applications of the functional equation for the family, as we will explain in Section 5. (Incidentally, the root number of the family is conveniently fixed in our theorem.) Nevertheless, the admissible class of test functions $\mathcal{C}_{\eta}$ is sufficiently large to deduce a version of Theorem 1.3 for incomplete $L$-functions.
Here is an extra benefit. The symmetries of all main terms for the cubic moment can now be addressed uniformly via archimedean Rankin–Selberg-type calculations in the style of Bump [Bump88] and Stade [St01]. The main result of [Bump88] only accounts for the diagonal term. To capture the symmetries of the remaining seven off-diagonal terms, we prove three new Mellin–Barnes identities.
@@
$J_{w,\alpha}(s;h)\$$=\ \frac{1}{4}\ \Gamma_{\mathbb{R}}(1+\alpha_{2}^{w}-\alpha_{3}^{w})^{-1}\Gamma_{\mathbb{R}}(1+\alpha_{1}^{w}-\alpha_{3}^{w})^{-1}$
$\hskip 40.0pt\ \cdot\ \int_{(0)}\,H(\mu)\,\prod_{\pm}\ \Gamma_{\mathbb{R}}(s\pm\mu+\alpha_{1}^{w})\Gamma_{\mathbb{R}}(s\pm\mu+\alpha_{2}^{w})\Gamma_{\mathbb{R}}(1-s\pm\mu-\alpha_{3}^{w})\ d^{W}\mu;$ (1.17)
@@
$\left(\mathcal{F}_{\alpha}H\right)\left(2s-1,s\right)\$= $=\ 2\,\prod_{i=1}^{3}\,\frac{\Gamma_{\mathbb{R}}(2s-1+\alpha_{i})}{\Gamma_{\mathbb{R}}(2-2s-\alpha_{i})}\,\int_{(0)}\,H(\mu)\,\prod_{i=1}^{3}\ \prod_{\pm}\,\Gamma_{\mathbb{R}}(1-s\pm\mu+\alpha_{i})\,d^{W}\mu$ (1.19)
for $w\in\left\{w_{2},w_{4},w_{\ell}\right\}$ and $i=1,2,3$.
@@
  • Cubic moment. See Proposition 5.1. The identities (1.17), (1.18) and (1.19) correspond to the “$1$-swap” terms (1.12), “22-swap” terms (1.13), and “$3$-swap” term (1.14), respectively. The diagonal term (1.11) is known as the “0-swap” term.
@@
    • The continuous spectrum: resulting in two “11-swap” terms and one “$2$-swap” term; see (1.16) and Lemma 6.3;
    • The non-constant degenerate part of the Fourier expansion: resulting in two “11-swap” terms and one “0-swap” term; see (1.15) and Proposition 6.10.
The notion of “swap” was introduced in [CK15a, CK15b] as a refinement of the CFKRS conjectures; see Sections 5.15.2.
Previously, Hughes and Young [HY10, Y11] investigated the main terms for the fourth moment of the $\zeta$-function by a method different from Motohashi’s. They employed the approximate functional equation (for the product of four $\zeta$’s) in conjunction with the $\delta$-symbol expansion (e.g., [DFI94]) applied to the shifted divisor sums, and careful smoothing and truncations. The upshot of their analysis is a set of exotic cancellations and combinations of many complicated-looking terms that, strikingly, reproduces the CFKRS predictions up to acceptable error. Petrow [Pe15, Sections 2-3] studied the full set of main terms for the cubic moment via an argument of comparable difficulty.
Our approach to extracting the full set of main terms differs from the existing ones. We are able to treat the main terms for both the cubic and fourth moments in a unified manner, without ad hoc cancellations or delicate combinations of terms. In Sections 6.8 and 7.1, we verify the agreement with the CFKRS conjectures by a short argument, using only Theorem 1.4, the relation $\alpha_{1}+\alpha_{2}+\alpha_{3}=0$, and the functional equation for the $\zeta$-function. Our perspective may shed light on recent developments in the mean value theorems for long Dirichlet polynomials [CK19, HN22, BTB22+, CF22+], which aim to identify the arithmetic sources of main terms in various families of moments of $L$-functions without the approximate functional equations.
For more abstract treatments of the main terms via regularized or degenerate functionals in automorphic representations; see [Ne20+, Section 11] and [Wu22, Section 1.5].
@@ 1.4.3. Integral transform
$(\mathcal{F}_{\alpha}H)(s_{0},1/2)\ =\$$2\,\Gamma_{\mathbb{R}}(1-s_{0})\prod_{i=1}^{3}\,\Gamma_{\mathbb{R}}(s_{0}+\alpha_{i})\int_{0}^{\infty}\,x^{1/2}\,\mathbf{F}\left(\begin{matrix}\frac{1-s_{0}}{2},\;\frac{1-s_{0}-\alpha_{1}}{2}\\ \frac{1}{2}\end{matrix}\Bigm|-x\right)\mathbf{F}\left(\begin{matrix}\frac{s_{0}+\alpha_{2}}{2},\;\frac{s_{0}+\alpha_{3}}{2}\\ \frac{1}{2}\end{matrix}\Bigm|-x\right)$
$\hskip 30.0pt\cdot\,\int_{(0)}\ H(\mu)\prod_{\pm}\,\prod_{\pm}\,\Gamma_{\mathbb{R}}(1/2\pm\mu\pm\alpha_{1})\mathbf{F}\left(\begin{matrix}\frac{1}{4}+\frac{\alpha_{1}+\mu}{2},\;\frac{1}{4}+\frac{\alpha_{1}-\mu}{2}\\ \frac{1}{2}\end{matrix}\Bigm|-x\right)\,\,d^{W}\mu\ d^{\times}x.$ (1.20)
@@ 1.5. Sketch, road map and discussion of our argument
Let $\mathrm{G}_{n}:=\mathrm{GL}_{n}(\mathbb{R})$, $\Gamma_{n}:=\mathrm{SL}_{n}(\mathbb{Z})$, $\mathrm{U}_{n}$ be the standard maximal unipotent subgroup of $\mathrm{G}_{n}$, $\mathbb{G}_{m}]:=\mathbb{Z}^{\times}\setminus\mathbb{R}^{\times}$, $\mathrm{U}_{n}]:=(\mathrm{U}_{n}\cap\Gamma_{n})\setminus\mathrm{U}_{n}$, and $\overline{\mathrm{G}}_{n}]:=\Gamma_{n}\setminus\mathrm{G}_{n}/\mathbb{R}_{>0}^{\times}$. In [Kw23], we study the analogue of Theorem 1.3 for a Hecke–Maass cusp form $\Phi:[\overline{\mathrm{G}}_{3}]\rightarrow\mathbb{C}$, and show that the corresponding Motohashi-type formula follows from a seemingly trivial equality for a “Fourier–Hecke period”:
@@
where $\Phi(g):=\Phi\begin{pmatrix}g&\\ &1\end{pmatrix}$  [^1] Some authors use the notation $(\mathbb{P}_{2}^{3}\Phi)(g)$ to indicate the projection (or restriction) of a function on $\mathrm{G}_{3}$ to $\mathrm{G}_{2}$. , and $\psi\in\mathrm{Hom}([\mathrm{U}_{2}],\,\mathbb{C}^{\times})$ is fixed and non-trivial. In [Kwa-thesis, Chapter 1.3.1], the overarching structure for implementing (1.21) is illustrated by a Reznikov diagram ([R08]):
@@
In this work, we investigate (1.21) in the case of the $\mathrm{SL}_{3}(\mathbb{Z})$ Eisenstein series associated with the isobaric sum $|\cdot|^{\alpha_{1}}\,\boxplus|\cdot|^{\alpha_{2}}\,\boxplus\,|\cdot|^{\alpha_{3}}$. Although periods of Eisenstein series play a fundamental role in many notable instances of spectral reciprocity, handling them rigorously remains highly nontrivial, and no general method is currently available. In addition to [Ne20+], we also mention [JN21+, Mi21+], which employ the regularization method of Ichino–Yamana [IY15], and [Za20+], which develops Michel–Venkatesh’s deformation and regularization techniques.
@@
In (1.21), the most immediate obstruction comes from the divergence of the integrals over $\mathbb{G}_{m}]$ when $\Phi$ is Eisenstein. Resolving this highlights the advantage of interpreting the classical Motohashi phenomena of (1.2) and Theorem 1.3, which are intrinsically dualities between $\mathrm{GL}(1)$ and $\mathrm{GL}(2)$ $L$-functions, through the periods and automorphic forms for the “larger” group $\mathrm{GL}(3)$. Our strategy differs from Nelson’s [Ne20+]. We crucially exploit the structure of the $\mathrm{GL}(3)\times\mathrm{GL}(2)$ Rankin–Selberg period, in which the unfolding does not rely on the full automorphy of the $\mathrm{GL}(3)$ form, at least within the region of absolute convergence (upon attaching the factor $|\det*|^{s-1/2}$ to (1.21) and taking $\operatorname{Re}s\gg 1$). This feature allows the spectral calculations in the cuspidal and Eisenstein cases to proceed in complete parallel, and it is worth noting that only the non-degenerate part of the Fourier–Whittaker expansion of $\Phi$ is used here.
@@
The development of the dual side of (1.21) embodies most of the novelties of this work and reveals several aspects not present in our earlier papers [Kw23, Kw25]. We begin by evaluating the unipotent period over $\mathrm{U}_{2}]$, which, upon invoking a suitable automorphy of $\Phi$ (Proposition 3.14), decomposes into a sum of three components (Proposition 6.6) via the Fourier–Whittaker expansion of $\Phi$ (Lemma 3.13). The unipotent integration eliminates the degenerate components of $\Phi$ responsible for divergence, by a two-line argument (Lemma 6.5). We are then in a position to evaluate the integral over $\mathbb{G}_{m}]$ for $\operatorname{Re}s\gg 1$ (again with the factor $|\det*|^{s-1/2}$). The remaining “benign” degenerate terms of $\Phi$ contribute to the main terms of our spectral reciprocity formula in non-trivial ways (Sections 6.36.4). We exploit the equivariance of the Fourier–Whittaker periods at several points, which offers significant advantages over the complicated Hecke combinatorics that typically arise from the $\mathrm{GL}(3)$ Voronoi formula in Kuznetsov-based approaches. Finally, by examining the explicit forms of each component in the development of (1.21) in terms of $\zeta$- and $L$-functions, we find that the whole expression admits analytic continuation to the region for $s$ described in Theorem 1.3. The rest of the main terms predicted by the CFKRS conjectures arise naturally as polar contributions in the continuation process. These are the subjects of Sections 6.5, 6.7 and 6.8.
@@
The unipotent integration also yields an ample supply of admissible test functions on the spectral side of our reciprocity formula, with their regularity and shapes tailored to standard ones in the literature of moments of $L$-functions. This is obtained through the Whittaker–Plancherel formula (Lemma 3.9) with a Poincaré series incorporated into the period construction (1.21) (Corollary 6.4). The Poincaré series further enables us to specify the archimedean component within the spectrum of $L^{2}([\overline{\mathrm{G}}_{2}])$, a feature particularly favourable for spectral applications, much as in the many uses of the Kuznetsov formula. The focus of this work is an exact, spectral version of Motohashi’s formula restricted to the spherical part of $L^{2}([\overline{\mathrm{G}}_{2}])$, complementing the results of [Mo93, BM05]; accordingly, we work with a spherical Poincaré series. If desired, one may similarly localize to another subspace of $L^{2}([\overline{\mathrm{G}}_{2}])$ by prescribing a different minimal $K$-type.
@@
One should be able to recover the results of [Mo93, BM05] with our method. In that setting, our analysis should be simpler—we do not require the Atkinson dissection (or any geometric dissection), and their four-variable continuation argument can be replaced by a single-variable one, as the “shifts” play no essential role in our analytic continuation argument (they are used only to simplify residual calculations). We leave these considerations for future work.
@@ 1.6. Additional Remarks
This type of mixed moments is also of interest to the literature due to the applications to simultaneous non-vanishing; for example, [BBFR20, Li09, Nu20+, X11].
We have not yet fully exploited the strength of our method in this work, particularly in its non-archimedean aspects. For example, in [Kw25], we establish a $\hbox{GL}(3)$ Motohashi-type formula that dualizes $\hbox{GL}(2)$ twists of Hecke eigenvalues into $\hbox{GL}(1)$ twists by Dirichlet characters. In the isobaric case, this yields a short proof of the key spectral identity of [Y11] and [BHKM20] for the fourth moment of Dirichlet $L$-functions.

### v3

@@ 1.4.2. Main terms
Recent works on Moment Conjectures (e.g., [64] and [23]) have shown that it is both technically and aesthetically preferable to state and prove results in terms of complete (rather than incomplete) $L$-functions. Indeed, the complete set of the main terms of a moment of $L$-functions is expected to reflect the symmetries originating from successive applications of the functional equation for the family, as we will explain in Section 5. (Incidentally, the root number of the family is conveniently fixed in our theorem.) Nevertheless, the admissible class of test functions $\mathcal{C}_{\eta}$ is sufficiently large to deduce a version of Theorem 1.3 for incomplete $L$-functions.
Here is an extra benefit. The symmetries of all main terms for the cubic moment can now be addressed uniformly via archimedean Rankin–Selberg-type calculations in the style of Bump [14] and Stade [70]. The main result of [14] only accounts for the diagonal term. To capture the symmetries of the remaining seven off-diagonal terms, we prove three new Mellin–Barnes identities.
@@
$J_{w,\,-\alpha}(s;h)\$$=\ \frac{1}{4}\ \prod_{i=1}^{2}\,\Gamma_{\mathbb{R}}(1-\alpha_{i}^{w}+\alpha_{3}^{w})^{-1}\int_{(0)}\,H(\mu)\prod_{\pm}\,\Gamma_{\mathbb{R}}(1-s\pm\mu+\alpha_{3}^{w})\prod_{i=1}^{2}\,\Gamma_{\mathbb{R}}(s\pm\mu-\alpha_{i}^{w})\,d^{W}\mu;$ (1.14)
@@
$\left(\mathcal{F}_{\alpha}H\right)\left(2s-1,s\right)\$= $=\ 2\,\prod_{j=1}^{3}\,\frac{\Gamma_{\mathbb{R}}(2s-1+\alpha_{j})}{\Gamma_{\mathbb{R}}(2-2s-\alpha_{j})}\,\int_{(0)}\,H(\mu)\prod_{j=1}^{3}\prod_{\pm}\,\Gamma_{\mathbb{R}}(1-s\pm\mu+\alpha_{j})\,d^{W}\mu,$ (1.16)
for $i=1,2,3$, $w\in\left\{w_{2},w_{4},w_{\ell}\right\}$ (see (3.29)), and $\alpha_{i}^{w}$’s defined in (3.3.3).
@@
  • Cubic moment. See Proposition 5.1. The identities (1.14), (1.15) and (1.16) correspond to the “$1$-swap” terms (1.10), “22-swap” terms (1.11), and “$3$-swap” term (1.12), respectively. The diagonal term (1.9) is also known as the “0-swap” term.
@@
    • The continuous spectrum: resulting in two “11-swap” terms and one “$2$-swap” term that can be collected into $\mathcal{R}_{\alpha}(1-s;H)$ in (1.7); see Lemma 6.3;
    • The non-constant degenerate part of the Fourier expansion: resulting in two “11-swap” terms and one “0-swap” term that can be collected into $\mathcal{R}_{-\alpha}(s;H)$ in (1.7); see Proposition 6.11.
The notion of “swap” was introduced in [19, 20] as a refinement of the CFKRS conjectures; see Sections 5.15.2 for detailed discussions. As already hinted in (1.14), the symmetries of the Weyl group of $\mathrm{GL}(3)$ are essential in understanding the structures of the main terms.
Previously, Hughes and Young [35, 73] investigated the main terms for the fourth moment of the $\zeta$-function by a method different from Motohashi’s. They employed the approximate functional equation (for the product of four $\zeta$’s) in conjunction with the $\delta$-symbol expansion (e.g., [27]) applied to the shifted divisor sums, and careful smoothing and truncations. The upshot of their analysis is a set of exotic cancellations and combinations of many complicated-looking terms that, strikingly, reproduces the CFKRS predictions up to acceptable error. Petrow [64, Sections 2-3] studied the full set of main terms for the cubic moment via an argument of comparable difficulty.
Our approach to extracting the full set of main terms differs from the existing ones. We are able to treat the main terms for both the cubic and fourth moments in a unified manner, without ad hoc cancellations or delicate combinations of terms. In Sections 6.8 and 7.1, we verify the agreement with the CFKRS conjectures by a short argument, using only Theorem 1.4, the relation $\alpha_{1}+\alpha_{2}+\alpha_{3}=0$, and the functional equation for the $\zeta$-function. Our perspective may shed light on recent developments in the mean value theorems for long Dirichlet polynomials [21, 32, 4, 18], which aim to identify the arithmetic sources of main terms in various families of moments of $L$-functions without the approximate functional equations.
For more abstract treatments of the main terms via regularized or degenerate functionals in automorphic representations; see [62, Section 11] and [71, Section 1.5].
@@ 1.4.3. Integral transform
$(\mathcal{F}_{\alpha}H)(s_{0},1/2)\ =\$$4\,\Gamma_{\mathbb{R}}(1-s_{0})\prod_{i=1}^{3}\,\Gamma_{\mathbb{R}}(s_{0}+\alpha_{i})\int_{0}^{\infty}\,\mathbf{F}\left(\begin{matrix}\frac{1-s_{0}}{2},\;\frac{1-s_{0}-\alpha_{1}}{2}\\ \frac{1}{2}\end{matrix}\Bigm|-x^{2}\right)\mathbf{F}\left(\begin{matrix}\frac{s_{0}+\alpha_{2}}{2},\;\frac{s_{0}+\alpha_{3}}{2}\\ \frac{1}{2}\end{matrix}\Bigm|-x^{2}\right)$
$\hskip 30.0pt\cdot\,\int_{(0)}\ H(\mu)\prod_{\pm}\,\prod_{\pm}\,\Gamma_{\mathbb{R}}(1/2\pm\mu\pm\alpha_{1})\mathbf{F}\left(\begin{matrix}\frac{1}{4}+\frac{\alpha_{1}+\mu}{2},\;\frac{1}{4}+\frac{\alpha_{1}-\mu}{2}\\ \frac{1}{2}\end{matrix}\Bigm|-x^{2}\right)\,\,d^{W}\mu\ dx.$ (1.17)
@@ 1.5. Sketch, road map and discussion of our argument
Let $\mathrm{G}_{n}:=\mathrm{GL}_{n}(\mathbb{R})$, $\Gamma_{n}:=\mathrm{SL}_{n}(\mathbb{Z})$, $\mathrm{U}_{n}$ be the standard maximal unipotent subgroup of $\mathrm{G}_{n}$, $\mathbb{G}_{m}]:=\mathbb{Z}^{\times}\setminus\mathbb{R}^{\times}$, $\mathrm{U}_{n}]:=(\mathrm{U}_{n}\cap\Gamma_{n})\setminus\mathrm{U}_{n}$, and $\overline{\mathrm{G}}_{n}]:=\Gamma_{n}\setminus\mathrm{G}_{n}/\mathbb{R}_{>0}^{\times}$. In [49], we study the analogue of Theorem 1.3 for a Hecke–Maass cusp form $\Phi:[\overline{\mathrm{G}}_{3}]\rightarrow\mathbb{C}$, and show that the corresponding Motohashi-type formula follows from a seemingly trivial equality for a “Fourier–Hecke period”:
@@
where $\Phi(g):=\Phi\begin{pmatrix}g&\\ &1\end{pmatrix}$  [^2] Some authors use the notation $(\mathbb{P}_{2}^{3}\Phi)(g)$ to indicate the projection (or restriction) of a function on $\mathrm{G}_{3}$ to $\mathrm{G}_{2}$. , and $\psi\in\mathrm{Hom}([\mathrm{U}_{2}],\,\mathbb{C}^{\times})$ is fixed and non-trivial. In [48, Chapter 1.3.1], the overarching structure for implementing (1.18) is illustrated by a Reznikov diagram ([69]):
@@
In this work, we investigate (1.18) in the case of the $\mathrm{SL}_{3}(\mathbb{Z})$ Eisenstein series associated with the isobaric sum $|\cdot|^{\alpha_{1}}\,\boxplus|\cdot|^{\alpha_{2}}\,\boxplus\,|\cdot|^{\alpha_{3}}$. Although periods of Eisenstein series play a fundamental role in many notable instances of spectral reciprocity, handling them rigorously remains highly nontrivial, and no general method is currently available. In addition to [62], we also mention [45, 55], which employ the regularization method of Ichino–Yamana [38], and [75], which develops Michel–Venkatesh’s deformation and regularization techniques.
@@
In (1.18), the most immediate obstruction comes from the divergence of the integrals over $\mathbb{G}_{m}]$ when $\Phi$ is Eisenstein. Resolving this highlights the advantage of interpreting the classical Motohashi phenomena of (1.2) and Theorem 1.3, which are intrinsically dualities between $\mathrm{GL}(1)$ and $\mathrm{GL}(2)$ $L$-functions, through the periods and automorphic forms for the “larger” group $\mathrm{GL}(3)$. Our strategy differs from Nelson’s [62]. We crucially exploit the structure of the $\mathrm{GL}(3)\times\mathrm{GL}(2)$ Rankin–Selberg period, in which the unfolding does not rely on the full automorphy of the $\mathrm{GL}(3)$ form, at least within the region of absolute convergence (upon attaching the factor $|\det*|^{s-1/2}$ to (1.18) and taking $\operatorname{Re}s\gg 1$). This feature allows the spectral calculations in the cuspidal and Eisenstein cases to proceed in complete parallel, and it is worth noting that only the non-degenerate part of the Fourier–Whittaker expansion of $\Phi$ is used here.
@@
The development of the dual side of (1.18) embodies most of the novelties of this work and reveals several aspects not present in our earlier papers [49, 50]. We begin by evaluating the unipotent period over $\mathrm{U}_{2}]$, which, upon invoking a suitable automorphy of $\Phi$ (Proposition 3.14), decomposes into a sum of three components (Proposition 6.6) via the Fourier–Whittaker expansion of $\Phi$ (Lemma 3.13). The unipotent integration eliminates the degenerate components of $\Phi$ responsible for divergence, by a two-line argument (Lemma 6.5). We are then in a position to evaluate the integral over $\mathbb{G}_{m}]$ for $\operatorname{Re}s\gg 1$ (again with the factor $|\det*|^{s-1/2}$). The remaining “benign” degenerate terms of $\Phi$ contribute to the main terms of our spectral reciprocity formula in non-trivial ways (Sections 6.36.4). We exploit the equivariance of the Fourier–Whittaker periods at several points, which offers significant advantages over the complicated Hecke combinatorics that typically arise from the $\mathrm{GL}(3)$ Voronoi formula in Kuznetsov-based approaches. Finally, by examining the explicit forms of each component in the development of (1.18) in terms of $\zeta$- and $L$-functions, we find that the whole expression admits analytic continuation to the region for $s$ described in Theorem 1.3. The rest of the main terms predicted by the CFKRS conjectures arise naturally as polar contributions in the continuation process. These are the subjects of Sections 6.5, 6.7 and 6.8.
@@
The unipotent integration also yields an ample supply of admissible test functions on the spectral side of our reciprocity formula, with their regularity and shapes tailored to standard ones in the literature of moments of $L$-functions. This is obtained through the Whittaker–Plancherel formula (Lemma 3.9) with a Poincaré series incorporated into the period construction (1.18) (Corollary 6.4). The Poincaré series further enables us to specify the archimedean component within the spectrum of $L^{2}([\overline{\mathrm{G}}_{2}])$, a feature particularly favourable for spectral applications, much as in the many uses of the Kuznetsov formula. The focus of this work is an exact, spectral version of Motohashi’s formula restricted to the spherical part of $L^{2}([\overline{\mathrm{G}}_{2}])$, complementing the results of [59, 12]; accordingly, we work with a spherical Poincaré series. If desired, one may similarly localize to another subspace of $L^{2}([\overline{\mathrm{G}}_{2}])$ by prescribing a different minimal $K$-type.
@@
One should be able to recover the results of [59, 12] with our method. In that setting, our analysis should be simpler—we do not require the Atkinson dissection (or any geometric dissection), and their four-variable continuation argument can be replaced by a single-variable one, as the “shifts” play no essential role in our analytic continuation argument (they are used only to simplify residual calculations). We leave these considerations for future work.
@@ 1.6. Additional Remarks
This type of mixed moments is also of interest to the literature due to the applications to simultaneous non-vanishing; for example, [2, 51, 63, 72].
We have not yet fully exploited the strength of our method in this work, particularly in its non-archimedean aspects. For example, in [50], we establish a $\hbox{GL}(3)$ Motohashi-type formula that dualizes $\hbox{GL}(2)$ twists of Hecke eigenvalues into $\hbox{GL}(1)$ twists by Dirichlet characters. In the isobaric case, this yields a short proof of the key spectral identity of [73] and [7] for the fourth moment of Dirichlet $L$-functions.